solve each system using the substitution method
x-2y=7
3x-21=6y

Answer by london maths tutor(243)


x-2y=7
3x-21=6y
if x = 13 and y = 3
Equation 1: x - 2y = 7
(13) - 2(3) = 7 , hence shown
Equation 2: 3x - 21 = 6y
3x - 6y = 21
3(13) - 6(3) = 21 , hence shown
Therefore, answer: x = 13 and y = 3